OpenStudy (rayford):

What is the general form of the equation of a circle with a center at the origin and a radius of 11?

Parth (parthkohli):

We know that the center is (0,0) and the radius is 11.

Parth (parthkohli):

Now to use the standard for of an equation of a circle:\[(x - h)^2 + (y - k)^2 = r^2 \](h,k) is the centre; r is the radius.

Parth (parthkohli):

Substitute the values, and play with the circle.

Parth (parthkohli):

\[ h = 0\\ k = 0\\ r = 11\]Continue please.

OpenStudy (rayford):

(x-0)^2+(y-0)^2=11^2

OpenStudy (rayford):

?

Parth (parthkohli):

Yeah. \(\text{some kinda number} - 0 = \text{some kinda number}\)

Parth (parthkohli):

So remember that \((x - 0)^2 = x^2\)

Parth (parthkohli):

Same thing with \((y - 0)^2\).

OpenStudy (rayford):

ok.

OpenStudy (rayford):

so..uhm

OpenStudy (rayford):

now what?

Parth (parthkohli):

Simplify 11^2 too.

OpenStudy (rayford):

121?

OpenStudy (rayford):

so... x^2+y^2=121?

OpenStudy (rayford):

may someone answer another one for me?

OpenStudy (alexwee123):

sure i love problems :o

OpenStudy (rayford):

and you love the olympics i see!

OpenStudy (alexwee123):

yep US BASKETBALL TMW :D

OpenStudy (rayford):

im more of a track dude lol

OpenStudy (alexwee123):

lol nice did you see the 400m relay final?

OpenStudy (rayford):

no :( but i did see like everything else!

OpenStudy (alexwee123):

lol nice nice :D well what's your question?

OpenStudy (rayford):

oh sorry...

OpenStudy (rayford):

What is the center and radius of the circle with equation x2 + (y – 8)2 = 25?

OpenStudy (alexwee123):

well the standard equation for a circle is  where (h,k) is the center and r is the radius so the center would be at (0,8) and the radius is square root of 25 or 5